# Acoustic BPE for Speech Generation with Discrete Tokens

### v1

@@ 1 Introduction
Discretization further compresses redundant information and enables the direct application of algorithms from natural language processing (NLP) communities. With $k$-means clustering or vector quantization, these aforementioned SSL features can be utilized as pseudo text for speech generation purposes. GSLM [8] extracts SSL features from pretrained HuBERT which is then discretized using $k$-means clustering. These discrete tokens are utilized as inputs to Tacotron2 [9] for synthesizing mel-spectrograms. VQTTS [10] [11] replaces the mel-spectrogram with vq-wav2vec tokens and a 3-dimensional prosody feature to bridge the acoustic model and the vocoder, achieving competitive synthesis quality compared to its mel-spectrogram counterpart. Building upon VQTTS, TN-VQTTS [12] further improves speaker adaptation by incorporating timbre-normalized vector quantized acoustic features. AudioLM [13] trains a language model on tokens discretized by w2v-BERT and $k$-means clustering, which demonstrates the effectiveness of discrete audio tokens as textual units in language modeling. This observation is further substantiated by SPEAR-TTS [14], which leverages discrete audio tokens as pseudo-textual units in low-resource scenarios.
However, the existing approaches directly utilize audio tokens, which poses challenges in sequence modeling. This is primarily due to the typically lengthy token sequences and the reliance on the model to capture correlations between tokens. While some solutions have been proposed to mitigate this issue, such as removing sequential repetitions of units used in [8][13], these approaches suffer from corruptive encoding, making them unsuitable for speech generation tasks. In text language modeling, this problem is relieved by combining a segment of consecutive tokens according to a certain rule. For example, Byte-pair encoding (BPE) [15] dynamically creates subword units based on frequency, encoding morphological information.
@@ 2.2 Speech language model
Given a sequence, the language model can calculate the probability of this sequence or predict the probability distribution of the next token to perform tasks such as continuation. The speech language model (SLM) extends language modeling to discrete audio tokens. However, unlike previous approaches [8][13][20] that directly utilize discrete audio tokens, our speech language model leverages acoustic BPE. This approach inherits the advantages of BPE encoding, such as enhanced sequence modeling and improved syntax capturing abilities. We use a decoder-only Transformer similar to [8][13] and briefly explain the process of training and generating continuations.
@@
Generating continuation: The speech language model can generate continuations after a given prompt $\bm{p}=\{p_{1},..,p_{l}\}$ by autoregressively sampling the next token $x_{i+1}$:
$x_{i+1}\sim p(x_{i}|x_{<i},\bm{p},\theta)$ (3)

### v4

@@ 1 Introduction
Discretization further compresses redundant information and enables the direct application of algorithms from natural language processing (NLP) communities. With $k$-means clustering or vector quantization, these aforementioned SSL features can be utilized as pseudo text for speech generation purposes. GSLM [8] extracts SSL features from pretrained HuBERT which is then discretized using $k$-means clustering. These discrete tokens are utilized as inputs to Tacotron2 [9] for synthesizing mel-spectrograms. VQTTS [10, 11, 12] replaces the mel-spectrogram with vq-wav2vec tokens and a 3-dimensional prosody feature to bridge the acoustic model and the vocoder, achieving competitive synthesis quality compared to its mel-spectrogram counterpart. AudioLM [13] trains a language model on tokens discretized by w2v-BERT and $k$-means clustering, which demonstrates the effectiveness of discrete audio tokens as textual units in language modeling. This observation is further substantiated by SPEAR-TTS [14], which leverages discrete audio tokens as pseudo-textual units in low-resource scenarios.
However, the existing approaches directly utilize audio tokens, which poses challenges in sequence modeling. This is primarily due to the typically lengthy token sequences and the reliance on the model to capture correlations between tokens. While some solutions have been proposed to mitigate this issue, such as removing sequential repetitions of units used in [8, 13], these approaches suffer from corruptive encoding, making them unsuitable for speech generation tasks. In text language modeling, this problem is relieved by combining a segment of consecutive tokens according to a certain rule. For example, Byte-pair encoding (BPE) [15] dynamically creates subword units based on frequency, encoding morphological information.
@@ 2.2 Speech language model
Given a sequence, the language model can calculate the probability of this sequence or predict the probability distribution of the next token to perform tasks such as continuation. The speech language model (SLM) extends language modeling to discrete audio tokens. However, unlike previous approaches [8, 13, 20] that directly utilize discrete audio tokens, our speech language model leverages acoustic BPE. This approach inherits the advantages of BPE encoding, such as enhanced sequence modeling and improved syntax capturing abilities. We use a decoder-only Transformer similar to [lakhotia2021generativeborsos2023audiolm] and briefly explain the process of training and generating continuations.
@@
Generating continuation: The speech language model can generate continuations after a given prompt $\bm{p}=\{p_{1},..,p_{l}\}$ by autoregressively sampling the next token $x_{i}$:
$x_{i}\sim p(x_{i}|x_{<i},\bm{p},\theta)$ (3)
@@ 4 Conclusion
5 Acknowledgements
This work was supported by China NSFC Project (No. 92370206), Shanghai Municipal Science and Technology Major Project
(2021SHZDZX0102) and the Key Research and Development Program of Jiangsu Province, China (No.BE2022059).